Solve 5 / 7 + 1 / 3 Solve 5 / 7 + 1 / 3 Solve 5 / 7 + 1 / 3

Solve
   5 / 7 + 1 / 3
Solve
   5 / 7 + 1 / 3
Solve
   5 / 7 + 1 / 3

Grade:12

2 Answers

Priyanshu Singh
3 years ago
use tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . comuse tiger algebra . com
Shraddha Singh
2 years ago
Dear Student,
 
5/7 + 1/3
L.C.M. of 7 and 3 is 21
= (5×3) + (1×7) / 21
= 15+7/21
= 22/21
 
Hope it helps!
Thank you